### v2

# In Situ Training of Implicit Neural Compressors for Scientific Simulations via Sketch-Based Regularization

Abstract
Focusing on implicit neural representations, we present a novel in situ training protocol that employs limited memory buffers of full and sketched data samples, where the sketched data are leveraged to prevent catastrophic forgetting. The theoretical motivation for our use of sketching as a regularizer is presented via a simple Johnson-Lindenstrauss-informed result. While our methods may be of wider interest in the field of continual learning, we specifically target in situ neural compression using implicit neural representation-based hypernetworks. We evaluate our method on a variety of complex simulation data in two and three dimensions, over long time horizons, and across unstructured grids and non-Cartesian geometries. On these tasks, we show strong reconstruction performance at high compression rates. Most importantly, we demonstrate that sketching enables the presented in situ scheme to approximately match the performance of the equivalent offline method.
keywords:
Continual learning , Neural fields , Streaming data , Scientific machine learning , Unstructured data , Catastrophic forgetting
journal: Elsevier
[…]
1 Introduction
[…]
Figure 1: The compression approach of this study applies to all mesh/geometry types, including (uniform) structured (left), unstructured (middle), and curvilinear (right), as it uses the node locations without connectivity information.
Neural network-based compression methods can capture non-linear phenomena and sharp features, such as physical discontinuities, that pose difficulties for traditional linear dimensionality reduction techniques, such as those based on singular value decomposition (SVD). Several neural compression approaches exist, which can compress the data directly or learn a reduced dimensionality representation, but all require optimizing neural network weights under some training strategy. One such method, implicit neural representation (INR) [park2019deepsdf, genova2019learning, sitzmann2020implicit], is an implicit, continuous, and differentiable neural network model that, conveniently, only requires access to the space-time coordinates of points, circumventing the connectivity structure entirely. INR, with some specific recent developments, has shown outstanding accuracy in representing a diverse set of signals; see, e.g., [sitzmann2020implicit, zhang2021implicit, pan2023neural, han2023kd, sales2024implicit] or Figure˜2 wherein we illustrate sample compression results for INRs trained offline on two complex simulation datasets.
[…]
Figure 2: Compression results for INRs trained offline. See Section˜3 for details on the datasets.
Our interest is in extending these INR compression techniques to the in situ setting, meaning the scenario where data samples (or points, entries, etc.) arrive one-at-a-time from a process (e.g., simulation) and cannot all be stored due to memory constraints. We are careful to use the term in situ as opposed to online or streaming because our procedure will not be able to run indefinitely and will not end precisely when the data generation process ends.
Available data samples are stored in what we will refer to as the buffer, which in the extreme case is only large enough to store a single sample, though we will typically target the case when the buffer can hold the equivalent of a few dozen data samples. Altering the size of this buffer, along with varying how many samples arrive together or what information from the samples is stored, allows one to interpolate between the pure online and offline settings. Our notion of samples is general and need not be a single number; in a typical example, a data sample is the entire spatial field at a given time (a snapshot), arising from a time-dependent partial differential equation (PDE) simulation. Non-stationary training for traditional machine learning techniques has been widely studied, but often less so in the case of neural networks, where it is typically referred to as continual learning. Hypernetworks, [ha2017hypernetworks], which are neural network architectures that learn to generate weights for a target network, are perhaps the most successful and well-known method for continual learning. Appropriated from multi-task learning [caruana1997multitask], they allow for a natural extension to in situ learning by equating one or multiple time steps with a task.
In this study, we present a novel learning procedure based on randomized sketching for in situ training of accurate and highly compressed neural representations of scientific simulation data using an INR-based hypernetwork architecture. Importantly, our methods are mesh-agnostic, only requiring access to nodes, so they are applicable to arbitrarily complex geometries. We highlight the fact that while sketching has been used extensively for low-rank compression, to our knowledge, this is the first time it has been used for regularizing neural compression. The remainder of this section will cover the necessary background and related work on INR, hypernetworks, and compression. Section˜2 will detail our in situ training, and Section˜3 extensively investigates our method’s practical performance on a variety of relevant datasets.
1.1 Preliminaries and Related Work
We will denote vectors using lowercase bold font (e.g., $\bm{x}$) and matrices or operators using uppercase bold font (e.g., $\bm{X}$). We will specialize to the task of compressing PDE simulation data and consider a fixed, possibly unstructured or non-uniform, mesh represented as $\bm{X}\in\mathbb{R}^{n\times d}$ consisting of $n$ points in $d$ spatial dimensions (typically 2 or 3). Additionally, we are provided with a $c$-dimensional time-dependent signal (or vector-field) $\bm{u}:\mathbb{R}^{d+1}\to\mathbb{R}^{c}$ sampled at all mesh nodes for time steps $t=1,2,\ldots,T$, generated by the numerical PDE solver [^1] For simplicity of notation, we assume a uniform time grid with unit spacing, but the method extends straighforwardly to nonuniform time steps.. For $c>1$, we may think of the different output coordinates of $\bm{u}$ as channels; for example, $\bm{u}$ might represent the 3D velocity field of a fluid, in which case $c=3$.
Although our core methods are based on points $(\bm{x},t)\in\mathbb{R}^{d+1}$, for training, we will often consider a single sample to be a snapshot, i.e., $\bm{U}_{t}=\bm{u}(\bm{X},t)$, $\bm{U}_{t}\in\mathbb{R}^{n\times c}$. We then define the reconstruction of this snapshot, via a neural network $\bm{\Phi}:\mathbb{R}^{d+1}\to\mathbb{R}^{c}$ parameterized by $\bm{\theta}$, as $\bm{\widetilde{U}}_{t}:=\bm{\Phi}(\bm{X},t;\bm{\theta})$. The concatenation of these into the full $T\times n\times c$ dateset and its reconstruction are given as $\bm{U}=[\bm{U}_{1};\cdots;\bm{U}_{T}]$ and $\bm{\widetilde{U}}=[\bm{\widetilde{U}}_{1};\cdots;\bm{\widetilde{U}}_{T}]$, respectively.
Our method will be used to optimize the parameters $\bm{\theta}$ of the neural network, whereupon denoting the number of parameters as $|\bm{\theta}|$, we can define the compression rate as follows:
$\frac{T\times n\times c}{|\bm{\theta}|}.$ (1)
Note that this does not include the required storage for the mesh or network definition, or intermediate storage for the in situ learning scheme. The formula assumes consistent floating point representation of all numbers, but should be adjusted in the obvious way if, for example, some of the network weights are represented in different precisions.
Implicit neural representation (INR)
INR is a straightforward form of supervised deep learning to train a neural network $\bm{\Phi}$ to approximate a target signal $\bm{f}$, and is the building block framework upon which have been built popular extensions, such as PINNs [raissi2019physics]. The key concept is to build a neural network that views input as discrete samples of a continuous object, so that dimensionality does not increase with resolution. Typically, the network in question is of a simple, densely-connected design, and the signal is a function of spatial or spatio-temporal coordinates (for notational simplicity, in this subsection we assume a purely spatial domain and hence denote the coordinates by $\bm{x}$). To be more precise, we consider a signal $\bm{f}:\mathbb{R}^{d}\to\mathbb{R}^{c}$ that maps coordinates in $d$ dimensions to a vector of $c$ features. Examples include 2D image pixel coordinates mapping to RGB colors, or mesh coordinates and time mapping to the flow field of a numerical PDE solution. Section˜1.1 defines the simplest INR with $L$ feedforward layers and non-linear activation function $\sigma$, but one should note that more complicated architectural features, such as residual connections, can be included:
$\bm{x}_{0}=\bm{x},$
$\bm{x}_{\ell}:=\bm{\phi}_{\ell}(\bm{x}_{{\ell}-1})=\sigma\left(\bm{W}_{\ell}\bm{x}_{{\ell}-1}+\bm{b}_{\ell}\right),\quad(\forall\ell=1,\ldots,L-1),$
$\bm{\Phi}(\bm{x})=\bm{W}_{L}(\bm{\phi}_{L-1}\circ\bm{\phi}_{L-2}\circ\cdots\circ\bm{\phi}_{1})(\bm{x})+\bm{b}_{L},$ (2)
for some trainable weight matrices $\bm{W}_{\ell}$ and biases $\bm{b}_{\ell}$, which are collected into the set of network parameters $\bm{\theta}$. A dataset consists of coordinate-value pairs $\{(\bm{x}_{i},\bm{f}_{i}=\bm{f}(\bm{x}_{i}))\}$, which one uses to optimize the parameters $\bm{\theta}$ of $\bm{\Phi}$, such that $\mathcal{L}(\bm{\Phi}(\bm{x}_{i};\bm{\theta}),\bm{f}_{i})$ is minimized for some target loss $\mathcal{L}$. It is not uncommon for the loss to be of Sobolev type, i.e., including a term targeting the difference of the derivatives (either exact or approximate) of $\bm{f}$ and $\bm{\Phi}$ with respect to some or all of the coordinates.
The activation function $\sigma$ has been shown to play a significant role in the performance of INR. Most notably, the work of [sitzmann2020implicit] demonstrated that a sinusoidal activation with a specific random initialization of the network weights is a vastly superior option compared to ReLU, Tanh, or other popular alternatives, calling their framework a SIREN network. Extensions of the periodic activation function have been considered by, e.g., [saragadam2023wire, roddenberry2023implicit], which introduce and analyze complex wavelets. See the comprehensive review [essakine2024we] for further details on the current state of INR.
Hypernetworks
Hypernetworks are a simple but powerful concept in deep learning wherein one network, the hypernetwork, is used to generate weights for a target network. Originally developed by [ha2017hypernetworks] as a form of weight sharing for convolutional and recurrent models, they have since been integrated with a variety of deep learning pipelines. Typically, the hypernetwork is much smaller than the target network and maps from an encoding of a specific task; for example, language pairs in translation. Less complicated architectures map directly to the full vector of target weights, but for large target models, this can be inefficient. Other approaches generate weights in chunks or via scaling factors. For details on the extensive work surrounding hypernetworks, see the review [chauhan2024brief].
Of particular relevance to our work is that of [pan2023neural], which combines hypernetworks and INR for learning low-dimensional representations of large scientific datasets. We use a similar sine-based INR with residual connections and generate time-dependent weights in the hypernetwork. Compression is not the focus of their work, and all models are trained offline, unlike our focus on in situ training and compression. Further, many of the stated disadvantages in the aforementioned work stem from a comparison to the SVD, which does not require hyperparameter tuning or network optimization. However, when the signal of interest exhibits a large Kolmogorov $n$-width, as in compressible and advection-dominated flows, linear dimensionality reduction techniques, such as the SVD, cease to be effective, thus motivating the use of non-linear dimensionality reduction, e.g., methods based on neural networks [lee2020model]. We also find that most of the remaining disadvantages, such as improving parameter efficiency or reducing training time, are deployment tasks, which can be tackled in isolation from the core compression problem.
Non-stationary learning
The machine learning community has used many synonymous or related terms, such as online learning, incremental learning, and continual learning, to describe training on data with a non-stationary distribution. This setting stands in contrast to the usual offline regime of most deep learning, where one assumes unfettered access to a fixed training dataset, which is used to optimize a model that is then deployed or tested on other data. In reality, a spectrum exists between fully offline learning and fully online learning, where samples arrive once and one at a time. Possible variations include multiple samples available at one time or the ability to store some data offline. The key difficulty lies in avoiding catastrophic forgetting, i.e., maintaining good performance for samples seen earlier in the training process [kirkpatrick2017overcoming].
Hypernetworks are especially adept at tackling these non-stationary problems, as the work of [von2019continual] highlights. To address catastrophic forgetting, they introduce a regularization term that seeks to maintain the hypernetwork output on previously learned tasks while learning a new task. A copy of the model is saved after learning each new task, which is used to reproduce the output of the hypernetwork on previous tasks. This method yields exceptional performance and limited forgetting, but the maximum number of tasks never exceeds 100. In our context, a task is a time step, and 100 is far fewer than the number of time steps often encountered in PDE simulations.
Our regularization approach using sketched spatial snapshots is more similar to the experience replay or rehearsal techniques commonly seen in reinforcement learning [rolnick2019experience] and throughout continual learning [wang2024comprehensive], though the technical details are notably different.
Compression
Both lossless and lossy compression for scientific data have long been targets of investigation. We emphasize the fact that this is a very dynamic field with significant amounts of research, much of it beyond our scope.
A large portion of this effort has been focused on error-bounded schemes for highly accurate compression of structured data, but at relatively small compression rates. Well-known tools include ZFP [lindstrom2014zfp] and SZ3 [liang2023sz], along with the recently introduced MFZ [doherty2024mfz] that extends to non-uniform data. See the recent review [di2025survey] for further details on error-bounded methods.
As mentioned earlier, significant effort has been put towards linear low-rank approximation methods for compression, motivated by the best [^2] In any unitarily invariant norm. rank-$k$ approximation from the SVD [eckart1936approximation]. Deterministic methods are often too computationally expensive for large datasets, so many methods turn to randomized decompositions. This employs randomized sketching [halko2011finding] — similar to what we consider in this work — but directly for compression as opposed to indirectly through regularization. We highlight the single-pass randomized methods using the SVD [yu2017single] and interpolative decomposition (ID) [dunton2020pass, pacella2022task, li2025online] for their focus on in situ compression.
One of the more significant directions of neural compression research uses INR in implicit neural compression (INC). Here, compression is accomplished by simply requiring that the size of the network is smaller than the size of the data so that Eq.˜1 is greater than $1$. Other approaches to compression include autoencoders, with recent developments to enable arbitrarily structured meshes [doherty2024quadconv]. A mixed autoencoder INR approach is also developed in [pham2023autoencoding]. The work of [yang2023tinc] introduces a blocking tree-based method with parameter sharing, and multiscale representations have also been considered in [saragadam2022miner, tang2024ecnr].
Most INC research has been focused on images [strumpler2022implicit, dupont2021coin] or video [zhang2021implicit, chen2021nerv], and some for general signals [pistilli2022signal]. More relevant to our task are [lu2021compressive, sales2024implicit], which apply INC to three and four-dimensional volumetric data from scientific simulations. Importantly, these works consider the offline scenario, incurring large training times and often using the entire dataset for meta-learning in a pre-training stage. Most also incorporate some form of quantization to reduce the model parameters, and thus increase the compression rate further.
A hybrid approach is considered in [madireddy2021], where a convolutional neural network is used to correct JPEG-based compression via artifact removal. A pre-trained model is fine-tuned in situ, in part using a buffer of stored data, before it is deployed to compress, i.e., remove artifacts, in a full simulation. While their use of experience replay ideas is similar to ours, the fundamental process is significantly different, given our sketch-based approach. Additionally, similar to many offline neural compression methods, their work requires significant processing outside of the in situ learning stage.
The most similar work to our own is the knowledge distillation approach of [han2023kd] (KD-INR), which also uses INR, but not hypernetworks. The difficulty associated with in situ learning is avoided by breaking the problem down into individual offline learning tasks. In particular, a new time-independent INR is trained at every time step, and then these are all distilled into a single time-dependent INR using a final knowledge transfer training stage. This final stage requires an extra training step, relative to our method, after the simulation has concluded. Additionally, this approach uses an offline pre-processing step to normalize the data and does not consider multi-channel training, although it is unclear how this impacts performance. Without the offline normalization step, KD-INR is also an in situ method, but it incurs different costs from our own approach, due to saving different objects. We did not compare directly to KD-INR because the code and data were not publicly available, so this is left to future work.
2 In Situ Training
The in situ setting is difficult because it limits the amount of training data the network is exposed to, and the sequential nature of the process can cause catastrophic forgetting of previously learned information. We propose extending INC to the in situ mode via a simple procedure that may be of independent use, as it does not depend on the network architecture in any particular way. Our procedure essentially falls under the experience replay umbrella as a method to regularize the learning process, but our twist is to do so with sketched data samples. Doing so enables us to store data from more time steps, as the size of that data is much smaller.
Figure 3: Our in situ INC training approach relies on a sketched data buffer to avoid catastrophic forgetting.
To ensure that the data is used most effectively in this single-pass regime, we first assume that we have some amount of available working memory, large enough to hold multiple snapshots at once. This memory is used to implement a buffer which holds data extracted from the simulation (Fig.˜3). The first portion of this buffer holds full snapshots in a queue of size $T_{f}\ll T$ — in fact, our numerical experiments use $T_{f}=1$. Newly generated snapshots are pushed into the queue, replacing old ones as the maximum size is reached. The second portion of the buffer contains sketched snapshots; the overall size of this buffer is not large, but its length $T_{s}\leq T$ can be quite long since the dimension of each sketch is very small. As we shall explain in further in Section˜2.1, sketching refers to a linear dimensionality reduction technique wherein a linear operator $\bm{S}:\mathbb{R}^{n\times c}\to\mathbb{R}^{k\times c}$ projects an $n$-dimensional vector onto a $k$-dimensional subspace, with $k\ll n$, such that some information in the original data is preserved. The total size of the stored buffer data is then given as follows:
$(T_{f}\times n+T_{s}\times k)\times c.$
The optimization loop then extracts a mini-batch (sub-sample) of full and sketched samples from the buffer to use in its update step. One should note that the simulation and optimization processes may be asynchronous, and in practice, many optimization iterations may take place in between new snapshots being generated.
To frame our in situ loss, it helps to consider the ideal loss we would use in the streaming case if one had no memory constraints. Let $\mathcal{L}$ be a given loss function, then the ideal loss for a reconstructed snapshot $\bm{\widetilde{U}}_{t^{\prime}}=\bm{\Phi}(\bm{X},t^{\prime};\bm{\theta})$ is given as
$\mathcal{L}_{\text{ideal}}(\bm{\Phi},t)=\frac{1}{t}\sum_{t^{\prime}=1}^{t}\mathcal{L}(\bm{U}_{t^{\prime}},\bm{\widetilde{U}}_{t^{\prime}}),$ (3)
which requires storing $\{\bm{U}_{t^{\prime}}\}_{t^{\prime}=1}^{t}$ for $t=1,\ldots,T$ and is thus impractical. Instead, we use Eq.˜4, for full and sketched batch sizes $b_{f}$ and $b_{s}$, respectively.
$\mathcal{L}_{\text{insitu}}(\bm{\Phi},t)=\underbrace{\frac{1}{b_{f}}\sum_{m=1}^{b_{f}}\mathcal{L}(\bm{U}_{t_{m}},\bm{\widetilde{U}}_{t_{m}})}_{\mathcal{L}_{\text{full}}}+\underbrace{\frac{\lambda}{b_{s}}\sum_{m=1}^{b_{s}}\mathcal{L}(\bm{S}_{t_{m}}\bm{U}_{t_{m}},\bm{S}_{t_{m}}\bm{\widetilde{U}}_{t_{m}})}_{\mathcal{L}_{\text{sketch}}},$ (4)
where $\lambda>0$ weights the sketch regularization term. In the case of $T_{f}=1$ the first term simplifies significantly. The mini-batch samples are always drawn anew at every iteration of the optimization algorithm, but the sketch $\bm{S}_{t^{\prime}}$ is only drawn once per time-step and then reused. It is also important to note that the product $\bm{S}_{t_{m}}\bm{U}_{t_{m}}$ is stored in the sketched buffer, as it cannot be reformed since we may not have saved the full snapshot. The sketched reconstruction $\bm{S}_{t_{m}}\bm{\widetilde{U}}_{t_{m}}$ is formed by sketching the full INR reconstruction, because only subsampling sketches can be applied by evaluating the network at a specific set (the subsample) of points.
Data: Sketch size $k$, buffer sizes $T_{f}$ & $T_{s}$, batch sizes $b_{f}\leq T_{f}$ & $b_{s}\leq T_{s}$
$\bm{\Phi}\leftarrow\text{build\_model}(~)$ ;
// Construct INC neural network
$\bm{B}_{f}\leftarrow\text{queue}(T_{f})$ ;
// Full queue
$\bm{B}_{s}\leftarrow\text{queue}(T_{s})$ ;
// Sketched queue
while simulation is running do
 case new snapshot $\bm{U}_{t}$ is available do
    $\bm{S}_{t}\leftarrow\text{construct\_sketch}(k)$ ;
    // Construct sketch
    $\text{push}(\bm{B}_{f},\bm{U}_{t})$ ;
    // Push full snapshot into queue
    $\text{push}(\bm{B}_{s},\bm{S}_{t}\bm{U}_{t})$ ;
    // Push sketched snapshot into queue
 for max training cycles do
    $\bm{U}_{f}\leftarrow\text{minibatch}(\bm{B}_{f},b_{f})$ ;
    // Sub-sample batch from full queue
    $\bm{U}_{s}\leftarrow\text{minibatch}(\bm{B}_{s},b_{s})$ ;
    // Sub-sample batch from sketched queue
    $\text{update\_model}(\bm{\Phi};\bm{U}_{f},\bm{U}_{s})$ ;
    // Optimizer step to update parameters
Algorithm 1 In Situ Training with Sketching
2.1 Sketching
Sketches may be deterministic or randomized. Often, in the latter case, we use random matrix ensembles based on specific distributions such as the Gaussian. Random subsampling can also be cast in this form by choosing a random subset of the $n$-dimensional standard basis. Gaussian ensembles generally perform well and have a plethora of theoretical results, but require slow, dense matrix operations. On the other hand, subsampling is extremely fast but can perform quite poorly in specific scenarios. In the worst-case, subsampling may select points that are the least helpful for the desired task, whereas dense sketches always aggregate information from all points. However, subsampling does preserve part of the mesh structure, as it is coarsening instead of aggregating, which can be beneficial in certain settings. Deterministic sketches with this preservation property have been studied in [dunton2021deterministic], with the idea that sketching operators informed by the data itself can outperform, both in speed and approximation quality, randomized sketches which are usually data oblivious.
It is important to make note of two aspects of a practical application of sketching. First, it is likely that the data or signal $\bm{u}$ has more than one channel, in which case the same sketch/transform is broadcast across them. Second, when the sketch is random, a random seed is generated for each snapshot and saved in the buffer alongside the sketched snapshot itself. When sketching for the loss in Eq.˜4, this seed is used in the random number generation for that sketch to ensure it is the same as the one used originally for the data. This does incur a minor amount of extra memory overhead due to the $T_{s}$ seed integers that must be stored alongside the sketched samples themselves, but this is minimal and avoids storing the entire sketch itself.
We elect to use the fast Johnson-Lindenstrauss transform (FJLT) [ailon2009fast] due to its superior speed and empirical performance compared to other possible options. However, we will also investigate the performance of subsampling in Section˜3 — and it is also used as a subroutine for the FJLT sketch — so we present this in Algorithm˜2. The FJLT is a Johnson-Lindenstrauss transform (JLT): a transformation that projects a finite set of points into many fewer dimensions while approximately preserving pairwise relative distances. The JL lemma (Theorem˜1) proves the existence of such a mapping, and any such mapping that satisfies this lemma may be referred to as a JL transform.
Theorem 1 (Johnson-Lindenstrauss Transformations [johnson1984extensions]).
For any $\epsilon>0$ and $n\in\mathbb{N}$, there exists a probability distribution $\mathcal{D}$ with support in $\mathbb{R}^{k\times n}$, where $k\leq\min\{n,\mathcal{O}(\epsilon^{-2}\log(m))\}$, such that for any set of $m$ points $\{\bm{p}\}_{i=1}^{m}\subset\mathbb{R}^{n}$, there exists $1>\delta>0$ such that if $\bm{S}\sim\mathcal{D}$ then the following holds with probability at least $1-\delta$:
$(1-\epsilon)\|\bm{p}_{i}-\bm{p}_{j}\|_{2}^{2}\leq\|\bm{S}\bm{p}_{i}-\bm{S}\bm{p}_{j}\|_{2}^{2}\leq(1+\epsilon)\|\bm{p}_{i}-\bm{p}_{j}\|_{2}^{2}\quad\forall i,j\in[m].$ (5)
In particular, for sufficiently large dimensions, using the probabilistic method, we see that there exists a linear map $\bm{S}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ with $k=\mathcal{O}(\epsilon^{-2}\log(m))$ such that Eq.˜5 holds.
Definition 1 (JL Transform).
If $\mathcal{D}$ is a probability distribution that satisfies the statements of Theorem˜1, we say $\mathcal{D}$ is an $(\epsilon,\delta)$-JL transform (with dimensions and sizes inferred from context). Following conventions in the literature, we often informally refer to a random matrix $\bm{S}\sim\mathcal{D}$ as a JL transform as well.
Furthermore, the construction of the distribution $\mathcal{D}$ is explicit, with the JLT first implemented via an orthonormalized Gaussian random matrix, for which Eq.˜5 holds with $k=\mathcal{O}(\epsilon^{-2}\log(m))$ [johnson1984extensions]. The practical downside is the slow, dense matrix-vector multiplications. The FJLT, on the other hand, uses a significantly different approach to gain a similarly significant speedup. Algorithm˜3 describes this process, which uses the orthogonal discrete cosine transform (DCT) to effectively mix rows [avron2010blendenpik] [^3] Specifically, a DCT of type II or III.. Algorithm˜3 satisfies Eq.˜5 with $k=\mathcal{O}(\epsilon^{-2}\log(m(\delta-n^{-\log^{3}(n)})^{-1})\log^{4}(n))$ (see, e.g., [huang2021spectral]) and, due to the DCT, only incurs an $\mathcal{O}(n\log(n))$ time cost since the DCT can be applied using fast Fourier transform techniques. As noted earlier, the random seed for the Rademacher random variable and subsampling are saved so that the same FJLT transform can be reused later.
Data: Data $\bm{U}\in\mathbb{R}^{n\times c}$, projected dimension $k\leq n$
$\Omega\leftarrow\operatorname{random\_choice}(\{1,\ldots,n\},k)$ ;
// Select $k$ indices uniformly without replacement
return $\bm{U}_{\Omega,:}$ ;
// Sample rows
Algorithm 2 Subsampling-Sketch $\bm{S}_{\text{subsample}}\bm{U}$
Data: Data $\bm{U}\in\mathbb{R}^{n\times c}$, projected dimension $k\leq n$
$\bm{d}\leftarrow\operatorname{rademacher}(n)\in\{\pm 1\}^{n}$ ;
// Independent, uniformly random signs
for $i\leftarrow 1$ to $c$ do
 $\widehat{\bm{U}}_{:,c}\leftarrow\sqrt{n/k}\operatorname{DCT}(\bm{U}_{:,c}\odot\bm{d})$ ;
 // $\odot$ is element-wise/Hadamard multiplication
return $\operatorname{Subsampling-Sketch}(\widehat{\bm{U}},k)$ ;
Algorithm 3 FJLT-Sketch $\bm{S}_{\text{FJLT}}\bm{U}$
2.2 JL Regularization for In Situ Learning
The goal of including the sketch loss in Eq.˜4 is to prevent forgetting past information from previously encountered samples and thus losing reconstruction accuracy. Section˜2.1 suggests informally that for a sketch satisfying Theorem˜1, we might be able to preserve the information embedded in the sketched data. But why should a sketch-based loss effectively serve as a regularization against catastrophic forgetting? In our main theoretical result, Theorem˜3, we will see that a sketch-based loss can serve as a surrogate to the true loss, i.e., that evaluated on the full data, for past samples.
First, we should clarify, for any given sketch, what exactly is being sketched. At first glance, for snapshot $t$, this is easily identified as the true data $\bm{U}_{t}$ and the reconstructed output of the INR on the entire mesh $\bm{\widetilde{U}}_{t}=\bm{\Phi}(\bm{X},t)$. However, this ignores the dependence of the INR on the parameters $\bm{\theta}$, which means we are sketching $\bm{\Phi}(\bm{X},t;\bm{\theta})$ for any possible $\bm{\theta}$. This violates Theorem˜1 because we are no longer sketching a finite set of points, given that the parameters are continuous. However, as we see in Theorem˜2, one can extend the Johnson-Lindenstrauss result to sketches on manifolds with an infinite number of points.
We refer to [Lee_manifoldBook2ndEd] for an extensive treatment of manifolds, but it is sufficient for our purposes to understand that a Riemannian manifold can be viewed as a subset of Euclidean space and behaves locally like a vector space. Informally, we are considering (possibly) lower-dimensional geometric structures embedded in some ambient subset of $\mathbb{R}^{N}$.
Theorem 2 (Manifold JL Transform [baraniuk2009random]).
Let $\mathcal{M}$ be a compact $M$-dimensional Riemannian manifold of $\mathbb{R}^{N}$ having volume $V$, condition number $\tau^{-1}$, and geodesic covering regularity $R$. Fix $\epsilon,\delta\in(0,1)$ and let $\bm{S}$ be a normalized random ortho-projector [^4] A random $k\times N$ matrix with orthogonal rows normalized by $\sqrt{N/k}$. from $\mathbb{R}^{N}$ to $\mathbb{R}^{k}$ which satisfies the following:
$k=\mathcal{O}\left(\frac{M\log\left(NVR\tau^{-1}\epsilon^{-1}\right)\log\left(\delta^{-1}\right)}{\epsilon^{2}}\right).$
If $k\leq N$, then with probability at least $1-\delta$ the following holds for $\bm{p},\bm{q}\in\mathcal{M}$:
$(1-\epsilon)\|\bm{p}-\bm{q}\|_{2}\leq\|\bm{Sp}-\bm{Sq}\|_{2}\leq(1+\epsilon)\|\bm{p}-\bm{q}\|_{2}.$
The volume, condition number, and geodesic covering regularity (the latter two of which are defined in [baraniuk2009random]) together define the topological regularity of the manifold. Note that this result, as originally stated, includes a factor of $\sqrt{k/N}$ as the sketch is not assumed to be normalized. Theorem˜2 presents a very similar result to Theorem˜1, but where the size of the sketch depends on geometric properties of the manifold, most importantly the intrinsic dimension, and not on the collection of points being sketched. This means that for low-dimensional manifolds of a potentially high-dimensional ambient space, one can construct a small JL transform for an arbitrarily large number of points. The size $k$ depends only logarithmically on the ambient dimension $N$, and in fact, this dependence can be removed completely as shown in subsequent works [ArminMike2015]. This ability to construct JL transformations for manifolds is key to our main result, which we present next.
Theorem 3 (JL Surrogate).
Define $\mathcal{L}$ as the squared $\ell_{2}$ (i.e., unnormalized MSE) loss. For a fixed mesh $\bm{X}$ and time $t\in[T]$, with parameters $\bm{\theta}\in\mathbb{R}^{N}$, suppose the map $\bm{\Phi}:\mathbb{R}^{N}\to\mathcal{M}\subset\mathbb{R}^{n\times c}$, defined via the neural network as $\bm{\Phi}(\bm{\theta})=\bm{\Phi}(\bm{X},t;\bm{\theta})$, maps to the Riemannian manifold $\mathcal{M}$. Further, suppose the sketch $\bm{S}_{t}:\mathbb{R}^{n\times c}\to\mathbb{R}^{k\times c}$ is a $(\epsilon,\delta)$-JL transform on the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, then, $\forall\bm{\theta}\in\mathbb{R}^{N}$, the following holds with probability at least $1-\delta$:
$\mathcal{L}(\bm{U}_{t},\bm{\Phi}(\bm{\theta}))\leq\frac{1}{1-\epsilon}\mathcal{L}(\bm{S}\bm{U}_{t},\bm{S}\bm{\Phi}(\bm{\theta}));$
and, also with probability at least $1-\delta$,
$\mathcal{L}(\bm{U}_{t},\bm{\Phi}(\bm{\theta}_{S}^{\star}))\leq\frac{1+\epsilon}{1-\epsilon}\min_{\bm{\theta}}\mathcal{L}(\bm{U}_{t},\bm{\Phi}(\bm{\theta})),$
where $\bm{\theta}_{S}^{\star}$ is a minimizer of the sketched loss.
Proof.
The unnormalized MSE loss is given as $\mathcal{L}(\bm{p},\bm{q})=\|\bm{p}-\bm{q}\|_{2}^{2}$. Via the left inequality of Theorem˜2,
$\mathcal{L}(\bm{U}_{t},\bm{\Phi}(\bm{\theta}))=\|\bm{U}_{t}-\bm{\Phi}(\bm{\theta})\|_{2}^{2}\leq\frac{1}{1-\epsilon}\|\bm{S}\left(\bm{U}_{t}-\bm{\Phi}(\bm{\theta})\right)\|_{2}^{2}=\frac{1}{1-\epsilon}\mathcal{L}(\bm{S_{t}}\bm{U}_{t},\bm{S}_{t}\bm{\Phi}(\bm{\theta})),$
which is the first desired inequality. Next, define $\bm{\theta}^{\star}$ as a minimizer of the full loss $\bm{\theta}\mapsto\mathcal{L}(\bm{U}_{t},\bm{\Phi}(\bm{\theta}))$ and $\bm{\theta}_{S}^{\star}$ as a minimizer of the sketched loss $\bm{\theta}\mapsto\mathcal{L}(\bm{S}\bm{U}_{t},\bm{S}\bm{\Phi}(\bm{\theta}))$. Then, we have the following:
$\mathcal{L}(\bm{U}_{t},\bm{\Phi}(\bm{\theta}_{S}^{\star}))$ $\leq\frac{1}{1-\epsilon}\mathcal{L}(\bm{S}\bm{U}_{t},\bm{S}\bm{\Phi}(\bm{\theta}_{S}^{\star}))$ via Theorem˜3
$\leq\frac{1}{1-\epsilon}\mathcal{L}(\bm{S}\bm{U}_{t},\bm{S}\bm{\Phi}(\bm{\theta}^{\star}))$ since $\bm{\theta}_{S}^{\star}$ is a minimizer
$\leq\frac{1+\epsilon}{1-\epsilon}\mathcal{L}(\bm{U}_{t},\bm{\Phi}(\bm{\theta}^{\star})),$ via right inequality of Theorem˜2
which is the second desired inequality. ∎
The proof of the second inequality is motivated by the informal proof of Theorem 2.14 in [woodruff2014sketching]. Also, note that we construct a JL transform for the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, which is just a convenience to simplify the proof, but does not introduce any complexity to the sketch itself. The shifted manifold will have the same geometric properties as $\mathcal{M}$ that impact the construction of the sketch.
The first bound in Theorem˜3 says that if we can find $\bm{\theta}$ to make the sketched loss small, then the true loss is also small, or, from another perspective, that for any given snapshot, the sketch loss serves as an approximate surrogate for the true loss. One may observe that the tightness of this bound is controlled by $0<\epsilon<1$, which also impacts the sketch size from Theorem˜2. For our goal of regularization, we may not need $\epsilon$ to be that small, thus allowing us to use smaller sketches. However, we don’t view this as a full explanation of our performance in practice. The trade-off that ideally holds, and which is empirically validated, is that much smaller sketches can still serve well enough as an approximate surrogate to the true loss to be an effective regularizer against forgetting. The actual learning is then performed based on the full snapshots.
The second bound says that we can make the minimizer of the sketched loss nearly as good as the minimizer of the true loss. We will make use of this bound in Section˜3 to loosely compare our chosen sketch sizes to what is suggested by the theory.
Theorem˜3 holds for a single snapshot, but in practice, we are sketching multiple snapshots at once to compute the overall sketch loss. The number of sketches actually depends on how long the process has been running and the size $T_{\text{sketch}}$ of the sketch buffer.
Theorem 4 (Batch JL Surrogate).
For a sketch buffer with size $T_{s}$, Theorem˜3 holds for the batch loss, i.e., the loss averaged over multiple snapshots, when the sketches are constructed with failure probability $\delta/T_{s}$.
Theorem˜4 extends the result of Theorem˜3 to multiple snapshots, which follows from a simple union bound over the individual sketches. To maintain the same probability of success, the probability of failure $\delta$ for each sketch must be decreased. This can be done at minimal cost due to the logarithmic dependence on this parameter in Theorem˜2.
We make a key assumption in Theorem˜3 that $\bm{\Phi}(\bm{\theta})$ maps to a Riemannian manifold, but is this actually true? Moreover, even if this assumption does hold, are the other relevant geometric properties from Theorem˜2 — dimension, volume, condition number, and geodesic covering regularity — in a way such as to yield a small sketch size? Rigorously answering this question is difficult. Considering just the dimension, we expect this to be dependent on both the data and the network parameters. For the former, in the worst case scenario, the data is entirely uncorrelated, so the dimension of $\mathcal{M}$ may be as large as $n\times c$. Similarly, in the best-case scenario, the data is perfectly correlated, so the dimension of $\mathcal{M}$ could be much smaller. In lieu of a theoretical answer to whether the INR output lies on a low-dimensional manifold, we provide computational evidence in Section˜3.2 that it indeed does.
3 Experiments
We conduct our experiments on the datasets listed in Table˜1, which highlight several potentially challenging or important attributes. The Ignition dataset describes a fully resolved gas wave front on a 2D uniform grid that transitions from a transport phase to a steady state jet. The Neuron dataset describes a 3D diffusion process in an unstructured, non-Cartesian branching neuron tree [Angran3D]. The Channel dataset, from the Johns Hopkins Turbulence Database [channel_flow_1, channel_flow_2], describes 3D turbulent flow on a non-uniform grid. It has been trimmed in space to a $64^{3}$ volume and in time to the first 500 snapshots. We define the notion of a “sample factor” as the relative size of the sketch to the mesh size for a given dataset, i.e., $100\cdot k/n$.
Mesh Type Spatial Dimension Snapshots Mesh Nodes Channels Memory (MB)
Datasets $d$ $T$ $n$ $c$
Ignition uniform 22 $450$ $2500$ 44 $18$
Neuron non-uniform 33 $500$ $116,943$ 11 $233.886$
Channel non-uniform 33 $500$ $262,144$ 33 $1572.864$
Table 1: Dataset descriptions. Total memory is calculated as $4\text{\times}{10}^{-6}$ MB per 32-bit floating point value.
Equation˜6 provides the frame-wise relative loss function we use for training, Eq.˜7 defines the full dataset relative Frobenius error, and Eq.˜8 defines the frame-wise peak signal to noise ratio (PSNR). The latter two of these three serve as our primary performance metrics.
$\mathcal{L}(\bm{U}_{t},\bm{\widetilde{U}}_{t})=\frac{1}{C}\sum_{c=1}^{C}\frac{\|(\bm{U}_{t})_{:,c}-\bm{(\widetilde{U}}_{t})_{:,c}\|_{2}}{\|(\bm{U}_{t})_{:,c}\|_{2}},$ (6)
$\text{RFE}(\bm{U},\bm{\widetilde{U}})=\frac{1}{C}\sum_{c=1}^{C}\frac{\|\bm{U}_{:,:,c}-\bm{\widetilde{U}}_{:,:,c}\|_{F}}{\|\bm{U}_{:,:,c}\|_{F}},$ (7)
$\text{PSNR}(\bm{U}_{t},\bm{\widetilde{U}}_{t})=\frac{1}{C}\sum_{c=1}^{C}20\cdot\log_{10}\left(\frac{\max_{i}(\bm{\widetilde{U}}_{t})_{i,c}}{\|(\bm{U}_{t})_{:,c}-(\bm{\widetilde{U}}_{t})_{:,c}\|_{2}}\right).$ (8)
Our experiments are designed to achieve two primary goals. First, to show that we can achieve strong performance, both offline and in situ, using our INC methods. Second, and more importantly, to show that we can approximately match the offline results using our in situ training paradigm. We do not optimize our models for maximum performance, nor do we further compress the data using other techniques such as network quantization. Further effort will likely yield improved results, but they are not the focus of this paper and are therefore left for future work. Instead, we choose simple models that yield desirable compression rates and relative error of a few percent.
Our work is implemented in PyTorch [pytorch] and available as an open-source GitHub repository [paper-repo]. All training is conducted on a single GPU (A100 or L40S) using the RAdam optimizer, a learning rate of $10^{-4}$, and in single precision (i.e., 32-bit floating point) [single_precision] for both the parameters and data. Given that our goal is to compress the given data, we do not seek to generalize to unseen data; thus, overfitting is not an issue. Our method is fully in situ, so we perform no pre-processing of the data. For all in situ results, the size of the full buffer is always $T_{f}=1$ (only one full snapshot available at a time), and the size of the sketch buffer is $T_{s}=T-1$ (one less than the full time horizon). We do not require sketching the final snapshot because the training loop ends after the last full snapshot has been encoded. The batch size of the sketch buffer and the number of epochs per full snapshot vary between datasets, but are in the range 25–45 and 300–500, respectively.
3.1 Model Architecture
As alluded to previously, we elect to use a hypernetwork approach for constructing our neural compressor. Practically, this results in two networks working together to map space-time coordinates to a reconstruction of the PDE solution field. Figure˜4 displays the generic INR architecture we employ for both the hypernetwork and the target network. The former maps from just the time coordinate $t$ to the parameters $\bm{\theta}(t)$. The latter ingests $\bm{\theta}(t)$ to define its layers and maps from the space-time coordinates $(\bm{x},t)$ to the reconstruction $\tilde{\bm{u}}(\bm{x},t)$.
(a) Network building-block consisting of two linear layers with sine activation functions and a skip connection.
(b) Overall network structure consisting of the hypernetwork and the target INR.
Figure 4: INC network structure.
Within the sine-layer blocks, we use sinusoidal activation functions with the initialization developed in [sitzmann2020implicit]. We also employ skip connections similar to those of [pan2023neural, sales2024implicit]. Thus, our networks are defined by two hyperparameters: the depth, i.e., the number of blocks, and the width of the internal linear layers. Because we are targeting the in situ setting, we cannot effectively normalize the data to ensure it is in some specific range, so no activation is used at the final linear layer. This ensures that the target network can produce any signal value. The lack of activation is also useful for the hypernetwork, as we construct the initial output of the hypernetwork to match the initialization of the target INR. To accomplish this, we use the strategy of [han2023kd], wherein the bias of the final linear layer in the hypernetwork is set to the exact values given by the initialization of the target INR, and the weights of the last linear layer are scaled by a relatively small factor.
We should note that many more tricks and techniques exist in the literature for both hypernetworks and implicit neural representation, much of which was discussed in Section˜1.1. Our architecture may benefit from these alterations. As an example, a more sophisticated hypernetwork architecture that does not map to all target parameters at once would be far more size-efficient, thus increasing our overall compression rate. However, this is left to future work.
3.2 Results
To put our proposed method, coupling Algorithm˜1 (buffer) with either Algorithm˜2 (subsampling) or Algorithm˜3 (FJLT), henceforth called “InSitu-FJLT” or “InSitu-Subsample” respectively, into context, we compare it with some baseline methods, both offline and in situ. We note that all the models use the same overall architecture but differ in how they are trained. Within each dataset, all models have the same number and size of SIREN blocks, and hence have the same compression rate. For offline methods, we first compare with a pure INC model that has access to all the data, using the loss $\mathcal{L}(\bm{\Phi},T)$ from Eq.˜3, and using mini-batches in time for training. This sketching-free method, which we label as “Offline-Baseline”, represents the best accuracy we could hope for using the INC framework. We also compare with “Offline-Subsample” and “Offline-FJLT”, which have access to all time points $t\in[T]$ but the spatial dimensions of $\bm{U}_{t}$ are sketched, either via Algorithm˜2 or Algorithm˜3, respectively. The subsampling sketch has no theoretical guarantees, but for the FJLT sketch, if the manifold hypotheses are valid and we choose sketches of size $k$ to give an $\epsilon$-JLT embedding, then we would expect these errors to be roughly a factor of $\frac{1+\epsilon}{1-\epsilon}$ worse than the full offline baseline. The “InSitu-Baseline” refers to using no sketched buffer at all (i.e., $T_{s}=0$), hence it is prone to catastrophic forgetting.
[…]
Offline-Baseline (full dataset) $2.17\pm 0.00$.19) $0.57\pm 0.00$.05) $5.37\pm 0.00$.36)
Offline-Subsample (sketched dataset, Algorithm˜2) $4.04\pm 0.00$.35) $14.1\pm 2.0$.4) $7.67\pm 0.00$.46)
Offline-FJLT (sketched dataset, Algorithm˜3) $3.87\pm 0.00$.49) $2.88\pm 0.02$.28) $7.30\pm 0.00$.26)
InSitu-Baseline (no sketched buffer) $84.2\pm 0.7$.2) $125.0\pm 2.3$.3) $77.2\pm 0.1$.6)
InSitu-Subsample, Algorithm˜1 + Algorithm˜2 $3.89\pm 0.02$.11) $21.0\pm 3.6$.3) $5.27\pm 0.00$.46)
InSitu-FJLT, Algorithm˜1 + Algorithm˜3 $2.64\pm 0.00$.85) $0.75\pm 0.00$.10) $5.22\pm 0.00$.37)
(a) RFE $\downarrow$ (%)
[…]
Offline-Baseline (full dataset) $41.6\pm 0.0$.6) $60.4\pm 0.0$.4) $37.4\pm 0.0$.4)
Offline-Subsample (sketched dataset, Algorithm˜2) $37.3\pm 0.1$.1) $35.4\pm 0.5$.6) $34.3\pm 0.0$.5)
Offline-FJLT (sketched dataset, Algorithm˜3) $37.1\pm 0.0$.9) $48.8\pm 0.6$.4) $34.6\pm 0.0$.2)
InSitu-Baseline (no sketched buffer) $9.28\pm 0.01$.84) $14.5\pm 0.1$.6) $14.3\pm 0.0$.44)
InSitu-Subsample, Algorithm˜1 + Algorithm˜2 $40.2\pm 0.1$.6) $40.7\pm 1.1$.8) $37.4\pm 0.0$.6)
InSitu-FJLT, Algorithm˜1 + Algorithm˜3 $41.7\pm 0.1$.2) $58.2\pm 0.0$.82) $37.4\pm 0.0$.5)
(b) PSNR $\uparrow$ (dB)
Table 2: Performance metrics on test datasets Ignition, Neuron, and Channel at 5%, 1%, and 2% sketch sample factors (where relevant), respectively. Metric statistics are reported as mean plus-or-minus standard deviation from five identical trials. The numbers in the parentheses of the first row of each table are compression ratios.
Table˜2 presents PSNR and RFE performance results on all three datasets in the offline and in situ settings. The key observation to be made is that sketching-based regularization often yields offline levels of performance in the in situ regime. In particular, sketching with the FJLT can yield results that approximately match those obtained offline across all datasets. While subsample sketching performs equally well on the Channel and almost as well on the Ignition datasets, its performance on the Neuron dataset is striking. Although it is not obvious from the presented statistics, the subsample results on the Neuron dataset are partially skewed by a few catastrophic failures. However, even in successful runs, the results are not at FJLT levels. Overall, we believe this result is not too surprising, as the non-Cartesian geometry of the neuron tree makes subsampling a particularly poor choice for sketching.
The results also show that our in situ methods, which combine training of the full snapshot with regularization from previous sketched snapshots, outperform the Offline-FJLT strategy (which just uses the sketched snapshots). The Offline-FJLT is conceptually simple to implement, since it does the sketching in situ and then trains offline. However, since it cannot train on the unsketched data, it incurs the $\frac{1+\epsilon}{1-\epsilon}$ error. Our in situ methods use sketched data but do not seem to incur this extra error, since the sketching is only used to prevent forgetting. This is also unsurprising, since for human memory, the task of learning something new is usually more difficult than the task of not forgetting something already learned.
In Table˜3, we explore how well the theory from Section˜2.2 aligns with our empirical results. Although unreported in Table˜2, we use Offline-FJLT and Offline-Baseline results for Eq.˜6 to compute the $\frac{1+\epsilon}{1-\epsilon}$ ratio from Theorem˜3, which we subsequently use to solve for $\epsilon$. We also compute principal component analysis (lPCA [fukunaga1971algorithm]) based estimates of the local manifold dimension for the INC networks used with each of the target datasets. This local dimensionality estimate yields a rough idea of the overall manifold dimension. We perform this analysis using the scikit-dimension package [scikit-dimension] by taking perturbations of size $10^{-5}$ around a nominal point of network parameters. This nominal state is obtained within the in situ setting after training on the first snapshot, i.e., the first time sketching is used.
Datasets Full Loss Sketch Loss Est. Manifold Dim. $M$ Est. Sample Factor (%)
Ignition $0.0262$ $0.0485$ $11$ $1.46$
Neuron $0.0076$ $0.0326$ $29.454$ $0.03$
Channel $0.0519$ $0.0717$ $50.008$ $0.19$
Table 3: Estimated sample factors using Theorem˜2 and Theorem˜3, with lPCA estimates of the local manifold dimension, for INC networks on the target datasets.
Combining the observed $\epsilon$ with the manifold dimension estimates, and ignoring log factors in Theorem˜2 (i.e., setting $k=M/\epsilon^{2}$), we compute the estimated sample factors. This is intended to be a rough estimate, but even with these approximations, it suggests that the Ignition dataset needs the greatest sketching dimension, followed by the Channel dataset, and finally the Neuron dataset. Our actual experiments used a $5\%$, $1\%$, and $2\%$ sample factor for the Ignition, Neuron, and Channel datasets, respectively, which matches the ordering suggested by the rough theoretical estimate. This suggests that the notion of manifold dimension of a dataset may indeed be a useful concept.
We can see visualizations of the results in Table˜2 comparing the original data to the offline baseline and in situ FJLT reconstructions for the Ignition and Channel datasets in Figs.˜5 and 6, respectively. For the Ignition dataset, Fig.˜5, we choose a snapshot from early in the simulation where the jet has not yet reached steady state. Note that the Ignition wavefront is accurately reconstructed by all methods, though the Offline Subsample method does have some small artifacts near the center of the wavefront, and the Offline FJLT has artifacts on the sides of the main jet. As observed in Table˜2, the in situ methods perform excellently, on par with the Offline-Baseline.
Figure 5: Reconstruction comparison of the Ignition data at snapshot $t=13$. The absolute error in the cutout is computed with respect to the original image, which is presented for reference. The colorbar only references the error cutouts.
For the Channel dataset, Fig.˜6, we select an arbitrary snapshot and $z$-coordinate for visualization, because the results do not differ across these dimensions due to the nature of the simulation. Across the three feature channels, the $x$-velocity (streamwise velocity) appears visually the worst. From a relative error perspective (RFE, Eq.˜7), however, this channel performs the best, at about 11%. The error of the other velocity channels, at around 686-8%, is what drags down the overall performance. We can attribute this paradox to the difference in scales between the velocity components and artifacts of the relative error metric. The streamwise velocity has a significantly larger scale than the other two components, so despite it appearing worse, it performs better relative to its scale.
Figure 6: Reconstruction comparison of the Channel dataset on all three channel flow velocities ($x,y,z$ components in top to bottom order) at snapshot $t=250$ for a constant $z$ coordinate in the center of the volume.
Per-snapshot relative error (Eq.˜6) curves for training and testing on the Ignition dataset are given in Fig.˜7. The training loss is split according to Eq.˜4 into the two components $\mathcal{L}_{\text{full}}$ and $\mathcal{L}_{\text{sketch}}$ corresponding to the two buffers. Thus, the “Full Loss” can be viewed as the instantaneous loss at the current snapshot, while the “Sketch Loss” is the averaged loss over previous sketched snapshots. The “Test Loss” is simply the instantaneous loss for a fixed snapshot after the network has been trained. For early snapshots, before approximately $t=50$, one can observe that the testing error is significantly better than the training loss, indicating the sketched snapshots have also actively been used for learning instead of only preventing forgetting.
Figure 7: Training and testing error on the Ignition data for the InSitu-FJLT case. Training error is taken at the end of the optimizer steps for a given snapshot. Data has been smoothed via convolution with a Gaussian.
Figure˜8 explores the impact of sketch size on reconstruction performance via five linearly spaced sample factors in $[0.001,0.01]$ and $[0.0055,0145]$ for FJLT and subsample sketching, respectively, on the Neuron dataset. In both RFE and PSNR metrics, FJLT performance improves consistently alongside the sample factor until reaching approximate offline values. As for subsampling, performance also increases with the sample factor, but we observe significantly more instability. In our experience, for larger sample factors, FJLT and subsample sketching lead to similar performance, but one may be ultimately limited in their choice by the amount of offline storage.
Figure 8: In situ Sketch sample factor performance comparison at fixed compression rate on Neuron dataset.
One may ask whether the inclusion of the hypernetwork in our overall INC architecture, shown in Fig.˜4, is necessary? Indeed, sketching as a regularization is just as applicable to a time-dependent INR, which we have found to be just as expressive in the offline setting (see Fig.˜2). To probe this question, Table˜4 presents comparison results on the Ignition dataset in the in situ regime with and without a hypernetwork. In the latter case, the hypernetwork branch seen in Fig.˜4(b) is simply eliminated. We include results for both our subsample and FJLT sketching options. Eliminating the hypernetwork results in fewer parameters, so the “Hypernet + INR” model is appropriately downsized to ensure approximately the same compression rate. We observe that the addition of a hypernetwork significantly increases performance regardless of the sketch.
Method PSNR $\uparrow$ (dB) RFE $\downarrow$ (%)
INR InSitu-Subsample $21.4\pm 0.1$.4) $49.6\pm 0.6$.7)
INR InSitu-FJLT $27.1\pm 0.1$.4) $22.6\pm 0.8$.1)
Hypernet + INR InSitu-Subsample $28.2\pm 0.9$.3) $32.0\pm 2.7$.5)
Hypernet + INR InSitu-FJLT $37.2\pm 0.0$.7) $4.7\pm 0.0$.5)
[…]
4 Discussion and Future Work
We have presented a new method for training neural compressors in situ with a simulation. The method primarily uses a limited memory buffer of sketched data samples to regularize against catastrophic forgetting during the learning process. This approach is theoretically motivated via a Johnson-Lindenstrauss-type result and empirically validated by showing that offline neural compression results can be matched in the in situ setting. In our particular case, we employ a hypernetwork paired with an implicit neural representation (INR), but the training approach can be extended to other network architectures. While our results are encouraging, a myriad of relevant questions remain unanswered that deserve further exploration. We break this into the two major components of our work: sketching for in situ learning and neural compression.
Sketching for In Situ Learning
Our investigation herein has only scratched the surface of how this paradigm could be used. To begin, the sketching techniques we considered — subsampling and the fast Johnson-Lindenstrauss transform (FJLT) — hardly comprise a complete overview of available options. Research on the best sketching method is warranted, including deterministic sketches or sketches that are data-informed in some regard.
Beyond the operator used to sketch, the protocol for when to sketch and how to incorporate them into the buffer has significant room for improvement. So far, we have considered identically sized sketches for each snapshot, but this is simply the most straightforward method to use the available memory. For example, it may be desirable to allocate more storage, allowing for larger sketches, to earlier snapshots. Additionally, an adaptive method for overwriting past sketch data would allow one to use all available memory at all times. As for the full snapshot buffer, we only considered the case where a single sample is available at a time. In practice, this number could be far larger, providing opportunities for sketching or encoding as a block.
On the theoretical side, much more can be done to explain the empirical success of small sketches in preventing catastrophic forgetting. The relevant question to ask is how large a sketch needs to be to serve as an effective regularizer? Rigorously answering this question would lead to a more precise setting of the sketch size. In a similar vein, our use of hypernetworks was empirically motivated, but deserves a more thorough examination as to why it works better with this regularization technique than a standard implicit neural representation (INR).
Neural Compression
Further work remains to make neural compression into a practical tool. While some improvements are straightforward, they have been set aside here to keep the focus on our primary contribution, i.e., demonstrating in situ training on an INR. For example, post-training (or other more complex approaches) network quantization can significantly increase the compression rate with little impact on the reconstruction performance. As well, we do not incorporate spatial gradients into our loss, but that is often an important aspect of sine-based INR (SIREN). On a similar note, we do not enforce any physics, despite having access to such knowledge from the simulation. Taking advantage of this extra information could improve the overall performance. In a more general sense, we also do not optimize the model architecture or its learning process.
In practice, many, if not most, large-scale simulations are carried out in a distributed manner, where the mesh is split into several similarly sized partitions. It is unclear how this change in setting would impact the neural compression approach. For example, one could optimize a central network via updates from the distributed partitions, or one could optimize distributed networks and then combine them in a distillation phase.
Acknowledgments
This material is based upon work supported by the Department of Energy, National Nuclear Security Administration under Award Number DE-NA0003968, as well as Department of Energy Advanced Scientific Computing Research Awards DE-SC0022283 and DE-SC0023346. Alireza Doostan’s work has also been partially supported by AFOSR grant FA9550-20-1-0138. We would also like to thank Kenneth Jansen, John Evans, Kevin Doherty, Angran Li, and Jeff Hadley from the University of Colorado Boulder for their helpful discussions surrounding this work. We also thank Youngkyu Kim for performing some insightful experiments regarding the performance of various sketches.
Additionally, this work was facilitated by several high-performance computing resources:
  • 1.
    The Alpine high-performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University, and with support from NSF grants OAC-2201538 and OAC-2322260.
  • 2.
    The Blanca condo computing resource at the University of Colorado Boulder. Blanca is jointly funded by computing users and the University of Colorado Boulder.
  • 3.
    Advanced computational, storage, and networking infrastructure provided by the Hyak supercomputer system and funded by the STF at the University of Washington.